Suppose a van de Graaff generator builds a negative static charge, and a grounded conductor is placed near enough to it so that a of negative charge arcs to the conductor. Calculate the number of electrons that are transferred.
Answer #1

Let q be the charge to the conductor. (Look up in your question as it is not mentioned here.)

Charge on one electron = 1.6*10^-19C

So, no. of electron n = q/(1.6*10^-19)

For example, if q = 9uC = 9*10^-6 C, then

n = (9*10^-6)/(1.6*10^-19)

= 5.63*10^13